Please help me solve this problem. The problem says to find the product and simplify:(4x-8)(x+2)
Answer by stanbon(75887)

find the product and simplify:(4x-8)(x+2)
Use FOIL:
F: 4x*x = 4x^2
O: 4x*2 = 8x
I: -8*x = -8x
L: -8*2 = -16
-------------------
Combine to get: 4x^2-16
===============================
Needless to say, this method is call FOIL.
===============================
Cheers,
Stan H.